### v2

# Unsupervised UAV 3D Trajectories Estimation with Sparse Point Clouds

Abstract
Compact UAV systems, while advancing delivery and surveillance, pose significant security challenges due to their small size, which hinders detection by traditional methods. This paper presents a cost-effective, unsupervised UAV detection method using spatial-temporal sequence processing to fuse multiple LiDAR scans for accurate UAV tracking in real-world scenarios. Our approach segments point clouds into foreground and background, analyzes spatial-temporal data, and employs a scoring mechanism to enhance detection accuracy. Tested on a public dataset, our solution placed 4th in the CVPR 2024 UG2+ Challenge, demonstrating its practical effectiveness. We plan to open-source all designs, code, and sample data for the research community @ github.com/lianghanfang/UnLiDAR-UAV-Est.
Index Terms—  Trajectory Estimation, UAV detection, Point Clouds, Unsupervised
1 Introduction
Drones have revolutionized various industries by enabling precise fertilizing in agriculture and allowing detailed inspection for hard-to-reach structures [1, 2]. However, the potential for malicious drone use is a significant concern. They can be exploited for unauthorized surveillance, drug trafficking, smuggling, and even deploying grenades in war zones. This threat highlights the urgent need for advanced detection systems to detect hostile drones effectively.
Detecting compact UAVs is challenging. Existing solutions rely on UAV control signals [3, 4] to detect, but can be bypassed easily by changing frequencies, using 5G networks, or fully autonomous drones [5, 6]. Visual-based methods [7, 8, 9] struggle with small objects at high altitudes. Narrow field-of-view cameras mounted on buildings can be manually operated to see the drone [10, 11, 12], but this is impractical for field operations. Wide field-of-view cameras can monitor a larger area but often only capture a few pixels of the drone as shown in Fig. 1 . Radar can detect drones effectively, but cheaper models are noisy [13] and expensive ones are costly and power-demanding [14]. Audio-based detection [15, 16, 17] are intuitive but often less effective, with most commercial drones being very quiet at a distance. LiDAR can detect drones, but its data is sparse at long ranges [18]. Overall, there is no perfect solution for drone detection.
Fig. 1: Illustration of detecting and tracking compact drones using a single low-cost sparse LiDAR to identify threats.
This work aims to accurately detect drones regardless of their control signal frequency or autonomy, including small drones at high altitudes, without manual operation. It ensures practicality for wide field operations and affordability for single-person or single-vehicle use, as shown in Fig. 1.
In this paper, We propose a concurrent clustering method for analyzing point clouds from a low-cost 3D LiDAR system. First, we perform global-local clustering to exclude large static objects. Then, we refine clustering using spatiotemporal density and voxel attributes to identify moving targets and isolate the UAV trajectory. Finally, we use spline fitting to reconstruct the UAV’s spatial trajectory, enhancing detection accuracy, reducing noise, and eliminating irrelevant data for clearer insights into drone movements.
Fig. 2: System Overview: Our algorithm uses DBSCAN to cluster point clouds, compares spatial-temporal changes, filters non-UAV data, and estimates UAV trajectories with spline fitting, measuring error with MSE.
The main contributions of our work are as follows:
  • Unsupervised Trajectory Estimation: We propose a fast, unsupervised method for detecting drone trajectories and positions from LiDAR point cloud data without any labels for supervised learning.
  • Spatio-Temporal Analysis: Our spatio-temporal voxel and density analysis method, with a scoring mechanism, isolates the correct trajectory point set.
  • Extensive Benchmarking: We benchmarked and tested various modalities with different methods to validate system performance. To our best knowledge, this is the first benchmark of its kind for this application.
  • Open-Source for All: We plan to open-source our design, codes, scripts and processed data for the benefit of the community and the general public github.com/lianghanfang/UnLiDAR-UAV-Est.
International Recognition: The proposed method is an improved iteration of our award-winning solution from the CVPR 2024 UG2+ Challenge, enhancing its cost-effectiveness, robustness, and reliability for practical field applications.
2 Related Works
This section reviews the limited literature on UAV detection and tracking, focusing on a few key approaches.
Vision-based UAV detection has evolved through deep learning, addressing challenges highlighted in studies such as Det-Fly [19] and MAV-VID, Drone-vs-Bird, and Anti-UAV [20]. Methods have improved accuracy by augmenting data and optimizing YOLOv4 for small UAV detection [21], and through transfer learning and adaptive fusion using simulated data [22].
Motion-assisted MAV detection integrates motion and appearance features using fixed and mobile cameras. Fixed camera methods employ background subtraction and CNN-based classification [23], while mobile cameras utilize spatiotemporal characteristics [24]-[25], but can struggle in dynamic environments. Another approach combines appearance and motion-based classification to distinguish MAVs from distractors [26], albeit facing challenges with similar moving objects.
Detection from moving cameras is complex due to the background and target motion mixing together. Methods using UAV-to-UAV datasets and hybrid classifiers [27] contend with background interference. Two-stage segmentation and feature super-resolution [28, 29] offer advancements but grapple with issues like motion blur and occlusions in complex settings.
LiDAR systems, widely used for object detection and tracking, face unique challenges with UAVs due to their small size, shape variability, diverse materials, high speeds, and unpredictable movements. One method adjusts LiDAR frame integration time based on drone speed and distance to enhance point cloud density and size, but this approach is intricate and sensitive to parameter settings [30]. Another strategy reduces LiDAR beams with probabilistic analysis and repositions the sensor for wider coverage, yet it struggles with continuous tracking of small points [31].
Fig. 3: This figure shows sampled points, ground truth, and our predicted trajectory, showing the accuracy of our solution.
Segmentation methods combined with object models and temporal information improve UAV detection and tracking effectiveness, though they are constrained by segmentation and object model accuracy [32]. Euclidean distance clustering and particle filtering algorithms offer accurate yet computationally efficient solutions, albeit sensitive to data noise and outliers [33]. In summary, while several methods address UAV detection and tracking challenges with LiDAR, each method presents distinct limitations and complexities, underscoring the need for ongoing research and development in this domain.
3 Proposed Methods
This section outlines our clustering-based, unsupervised spatial-temporal approach for detecting MAVs under challenging conditions. An overview of the system is shown in Fig. 2.
3.1 Global-Local Point Set clusterings
Let $\mathcal{F}$ denote a sequence of LiDAR scan frames, with $f$ denoting the number of frames. $\mathcal{P}$ represents a set of 3D points from a single scan in $\mathcal{F}$ The number of points in set $\mathcal{P}$ is denoted $card\,\mathcal{P}$. $\mathcal{C}$ denotes a cluster (subset) of points from $\mathcal{P}$ and $\rho$ denotes the density of points, $\mathcal{V}$ denotes the voxels of a set of points.
For local representation, $\left(\mathcal{P}|\mathcal{F}_{i}\right)$ represents the set of points $\mathcal{P}$ within the $i$-th frame $\mathcal{F}$. $\mathcal{C}_{j}\left(\mathcal{P}|\mathcal{F}_{i}\right)$ denotes the $j$-th cluster category of points from $\left(\mathcal{P}|\mathcal{F}_{i}\right)$. For global representation, $\sum_{i}^{j}\mathcal{F}$ denotes frame sequences from frame $i$ to frame $j$. $\mathcal{C}_{k}\left(\mathcal{P}|\sum_{i}^{j}\mathcal{F}\right)$ denotes the $k$-th category of the cluster after merging the points from $\sum_{i}^{j}\mathcal{F}$.
To distinguish between the results of global and local clustering, $\mathcal{C}_{k}\left(\mathcal{P},\mathcal{F}_{n}\mid\sum_{i}^{j}\mathcal{F}\right)$ represents the $k$-th cluster of points in the $n$-th frame, where the clustering is derived from the sequence of frames $\sum_{i}^{j}\mathcal{F}$.
And $\mathcal{V}\left(\mathcal{C}\right)$ indicates the size of the voxel occupied by cluster $\mathcal{C}$ in the space. Let $\rho_{\mathcal{C}}$ be an operator that denotes the density of points in a cluster $\mathcal{C}$ derived from a set of points $\mathcal{P}$ in the context of a frame sequence $\mathcal{F}$.
We first superimpose the point cloud on the global time frames to obtain $\left(\mathcal{P}|\sum_{i}^{j}\mathcal{F}\right)$, then use DBSCAN to perform clustering to obtain $\mathcal{C}\left(\mathcal{P}|\sum_{i}^{j}\mathcal{F}\right)$.
We calculate the density $\rho_{\mathcal{C}}(\mathcal{P}\mid\mathcal{F})$ of the point cloud in global point set.